OpenStudy (jamesreed30):

Find the area of a circle circumscribed about an equilateral triangle whose side is 18 inches long

OpenStudy (eliesaab):

Let x=18 be the side, then the height of the triangle is \[h=\frac{x \sqrt 3}{2} \]

OpenStudy (eliesaab):

The area is \[ \frac {x h }2 \]

OpenStudy (eliesaab):

Replace and you get your answer

OpenStudy (eliesaab):

\[ Area= \frac { x^2 \sqrt 3}4= \frac{18^2 \sqrt 3}4=81 \sqrt 3 \]

OpenStudy (jamesreed30):

So divide 81 and 3 and that's my answer?

OpenStudy (jamesreed30):

Also my answer has the PI symbol at the end so would that change by answer?

OpenStudy (jamesreed30):

@eliesaab

OpenStudy (eliesaab):

Oh Sorry, I found the are of the triangle. You need the one for the circle. Wait a minute

OpenStudy (jamesreed30):

okay

OpenStudy (eliesaab):

The radius of the circle is one third of the height h. So \[ Circle Area = \pi r^2= \pi (1/3 h)^2=\frac {h^2}9 \pi=\frac{\pi}9 h^2 \]

OpenStudy (eliesaab):

Now \[ h=\frac{x \sqrt 3}2\\ CircleArea=\frac \pi 9 \left( \frac{x \sqrt 3}2\right)^2 \] Replace x by 18 and you are done

OpenStudy (jamesreed30):

So 6 PI?

OpenStudy (eliesaab):

No

OpenStudy (eliesaab):

\[ CircleArea=\frac \pi 9 \frac { 3 x^2}4=\frac \pi {12} x^2=\frac \pi {12} 18^2= \frac \pi{12}324=27 \pi\]

OpenStudy (jamesreed30):

Thank you

OpenStudy (eliesaab):

Let me double check my computations

OpenStudy (eliesaab):

The answer is \[27 \pi \]

OpenStudy (jamesreed30):

Alright